OpenStudy (jmansedg):

5! = A. 14 B. 25 C. 120 D. 240

OpenStudy (jmansedg):

@silveralchemist09

OpenStudy (anonymous):

I would say b

OpenStudy (anonymous):

no direct answers

OpenStudy (jmansedg):

what is it silver well explain it

OpenStudy (anonymous):

C

RhondaSommer (rhondasommer):

the only way to really explain it is to put it in google calculator.

OpenStudy (anonymous):

5*4*3*2*1 = 120

OpenStudy (anonymous):

4r = 3s --> eqn 1, we'll have s = 4r/3 2t+10 = 3t - 15 Hence, t = 25 2t + 10 = 3t - 15 = 60 from summation angle of quadrilateral equals 360, we have (2t + 10) + (3t - 15) + 4r + 3s = 360 60 + 60 + 4r + 3(4r/3) = 360 120 + 8r = 360 8r = 240 r = 30

OpenStudy (jmansedg):

There are eight green shirts and four orange shirts hanging in a closet. If you randomly grab one shirt from the closet, is there a greater probability of grabbing a green shirt or an orange shirt? A. Green shirt B. Orange shirt

OpenStudy (anonymous):

B

OpenStudy (jmansedg):

16! = A. 136 B. 256 C. 57,657,600 D. 20,922,789,888,000 is it d

OpenStudy (anonymous):

the answe is c.

OpenStudy (jmansedg):

10! = A. 54 B. 1,814,400 C. 3,628,800 D. 7,557,600

OpenStudy (anonymous):

A. 54

OpenStudy (jmansedg):

If you are conducting a survey to determine which American cities are the most exciting in which to live, which of the following samples would be the least biased and most useful? A. Citizens from each city in the running B. Citizens from various cities all across America C. Group of foreign tourists visiting America D. Political leaders in each city in the running

OpenStudy (anonymous):

d.

OpenStudy (jmansedg):

ok i am turning it in right now letts see

OpenStudy (anonymous):

ok

OpenStudy (jmansedg):

i got an 70 it is passing so ya

OpenStudy (anonymous):

great you passed